OpenStudy (ryan0716):

negative 1 over 2 x + 1 = –x + 8

OpenStudy (anonymous):

x^(-1/2) = 5 + sq.rt. (1/12) (1)/(x)^1/2 = 5 + (1/12)^(1/2 (1)/(x)^1/2 = 25^1/2 + (1/12)^1/2

OpenStudy (anonymous):

medal and fan would be nice

OpenStudy (♪chibiterasu):

Um

OpenStudy (♪chibiterasu):

It's: \[-\frac{ 1 }{ 2 } x\] Not: \[x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

x=14

OpenStudy (ryan0716):

thx for the help everyone